# Dual-component stellar assembly histories in local elliptical galaxies via MUSE

### v1

@@ 3.1 Galaxy decomposition with BUDDI
Having modelled the surface brightness of the galaxies with one, two, and three Sérsic profiles, we opted for the two-component models for all of them, which are shown in Figure 1. To unveil the physical properties and implications associated with the different components of elliptical galaxies, the clean spectrum of each component was extracted with BUDDI. The magnitudes from GalfitM correspond to the total fluxes of each Sérsic profile integrated out to infinity. This extends to the individual image slice models in the datacube, and is reflected in the final 1D spectrum derived by BUDDI, which is obtained by plotting these fluxes as a function of wavelength (Johnston et al., 2017). The decomposed spectra of the inner and outer components, the sky, and that corresponding to the best fit model (inner + outer + sky) are shown for J020536 in Figure 2, and in Figures 10 and 11 for J205050 and J225546 respectively. We note that since most of the observing nights for the filler programme included high lunar illumination (Sect. 2), we have included an arbitrarily scaled scattered lunar model spectrum for each galaxy, obtained from the ESO SkyCalc 444Available from https://www.eso.org/observing/etc/bin/gen/form?INS.MODE=swspectr+INS.NAME=SKYCALC Sky Model Calculator tool (Noll et al., 2012; Jones et al., 2013). The effect of scattered moonlight is strong for galaxies J205050 and J225546, and closely matches the slope of the sky spectrum modelled by BUDDI, effectively demonstrating its ability to cleanly disentangle the light from physical galaxy components and from sky contributions. For J020536, the scattered moonlight has less of a contribution to the modelled sky spectrum. However, since the MUSE datacubes undergo two sky subtraction steps (once within the EsoRex pipeline, and later enhanced with the ZAP code), the lunar contribution could in principle be partially eliminated, leading to the slope differences between the scattered moonlight spectrum and the sky spectrum, especially in J020536. An important point to note is that wherever there is a known telluric feature in the lunar spectrum (for instance, the O2 A-band at Å in absorption), there appears a spectral feature in the sky spectrum modelled by BUDDI. The spectral feature in the sky spectrum may or not be in absorption due to the uncertainties in sky subtraction for the MUSE cube (including possible over-subtraction in some cases); it is nevertheless encouraging that BUDDI picks up these features at the expected wavelength. Software is often only as effective as the quality of the data it processes - even so, it has clearly demonstrated its effectiveness in cleanly decomposing the component spectra from the sky spectrum. The decomposed spectra of the components were then used for the stellar population analyses, in order to unveil the physical properties and implications associated with the different components of elliptical galaxies.
@@ 3.2 Spectral fitting for stellar population analysis
The spectra of the two components obtained from the previous step (Sect. 3.1) were fit using pPXF to derive the stellar metallicity [M/H] and stellar age. The Python implementation of pPXF v9.1.1 was used in this work. pPXF finds a linear combination of stellar templates that best matches the input spectrum. For this study, both the mass-weighted and light-weighted stellar population properties were derived and analysed. The full spectral fitting method optimises the mass or light weights for each SSP spectrum in the template library, such that the observed spectrum is a combination of these weights (Cappellari & Copin, 2003; Cappellari, 2017, 2023). The mean logarithmic mass and light-weighted stellar ages and metallicities estimated by pPXF are defined as:
@@
The empirical library of stellar spectra employed is MILES (Vazdekis et al., 2015)555Available from http://miles.iac.es/pages/webtools.php, along with the BaSTI isochrone models (Pietrinferni et al., 2004), which track stellar evolution. This choice of the stellar library is attributed to the vast coverage of stellar metallicities and ages, which is necessary for studying nearby galaxies in the local Universe, which can contain both old and young stars with low and high metallicities. Restricting to the safe ranges defined in Vazdekis et al. (2010), the chosen stellar population parameters span a range from -1.79 to 0.26 dex for the metallicities [M/H] and from 30 Myr to 14 Gyr for the ages, forming a combined total of 530 template spectra. Since each stellar template is normalised to 1, pPXF derives the mass weights by default. In order to estimate the light weights, the stellar templates are normalised by the band luminosity ( Å), chosen such that the resulting properties favour more recent star formation episodes in the galaxy. The Universal Kroupa IMF (Kroupa & Boily, 2002) with a slope of 1.3 is assumed.
@@
During the spectral fit with pPXF, an 8th-order multiplicative Legendre polynomial was used to correct for the shape of the continuum. Furthermore, the inclusion of a multiplicative polynomial in the fit also ensures that the spectral fit is independent of dust reddening, thereby eliminating the need to define a reddening curve explicitly within the pPXF routine (Cappellari, 2017). Additionally, the [OI]5577Å sky line is masked during the fit. The code allows fitting of emission lines simultaneously alongside the stellar continuum using a set of respective Gaussian templates representing the nebular emission. The spectral range for the fit is constrained to rest-frame Å due to inconsistencies that can be introduced when using the full MUSE wavelength range (Krajnović et al., 2015). Moreover, this chosen range spans the important spectral features that are sensitive to stellar age and metallicity (, , , and Fe) and aligns with the wavelength coverage in the MILES models.
@@
The pPXF routine provides a solution to the ill-posed problem of extracting stellar population properties from observed spectra. The extensive span of the parameter space of the SSPs inevitably leads to non-unique solutions that are degenerate in metallicity or age. This is one of the commonly cited challenges in the fossil-record method with stellar population modelling and full spectral fitting, along with the imprecise age resolution for older populations with Gyr (Cid Fernandes et al., 2005; Riffel et al., 2009; Ibarra-Medel et al., 2019; Lacerna et al., 2020; Riffel et al., 2024). The solution incorporated within pPXF is regularisation, which is responsible for mitigating intrinsic degeneracies in the final solution. When a possible degeneracy is encountered, the inclusion of a regularisation term treats this by smearing out the weights of similar metallicities and ages, going from discrete weights that resemble a single stellar population to a solution with smooth weights that are more physically motivated in galaxies. The REGUL parameter within the routine sets the degree of smearing, following the prescription outlined in the pPXF user manual and in Shetty & Cappellari (2015) and Jegatheesan et al. (2024). A word of caution is advised at this point while interpreting the star formation histories from full spectral fitting: star-formation is a stochastic and unpredictable process, which can occur at timescales shorter than those specified in the synthetic template models. The spectral fits for each component spectrum in the three elliptical galaxies resulting from pPXF are displayed in Figure 3. We find substantially robust fits for all objects, with the exception of the outer of component of J205050. From Figure 10, this galaxy appears to have the strongest lunar contamination, which has mostly been modelled off as the sky spectrum. However, if there is a fraction of residual scattered moonlight contaminating especially the outer regions, this can add to the noise of the spectrum of the outer component. This could lead to a less robust fit to the spectrum using pPXF; however, we still find that the crucial spectral features required for stellar population analysis have been modelled fairly well, and though we caution the direct interpretation of the fits and the resulting stellar populations for this component, we note that the star formation histories are consistent with the other galaxies.
@@ 3.3 Spectral index measurements for stellar population analysis
We additionally measure the line strengths of stellar age and metallicity sensitive absorption features in the spectra through the Lick indices defined in Worthey (1999). This analysis was done as a consistency check to empirically determine the stellar populations from the spectra, due to potential model-dependent uncertainties and the effects of trade-offs achieved with the choice of the regularisation parameter, that can be introduced in the age-metallicity distribution estimated through full spectral fitting with pPXF. The absorption line strengths were measured using the public Python library PyLick666Available from https://gitlab.com/mmoresco/pylick/ (Borghi et al., 2022), which performs a fast estimation of the predominant Lick indices and their uncertainties, spanning the UV to the near-infrared region. The uncertainties are estimated following the method of Cardiel et al. (1998), through simulations accounting for the errors in the LOS velocities and S/N in the observed spectrum.
The H index served as the spectral diagnostic for stellar age, while the metallicity was determined using the composite index [MgFe]′ = , which is less sensitive to the -element abundance and serves as a more robust indicator of the total metallicity of the galaxy or component. These indices were plotted on a H[MgFe]′ plane in Figure 5, overlaid with the SSP model grids defined in Vazdekis et al. (2010). The chosen grid spans ages from Gyr and metallicities [M/H] from -1.79 to +0.26 dex, constructed with the Lick index measurements from the MILES stellar library (Sánchez-Blázquez et al., 2006). The flexibility of the MILES webtools lies in its ability to customise the SSP models by convolving the model spectra with the velocity dispersion of the galaxy, effectively accounting for line broadening effects in the observed spectrum. As such, we use a convolved model with a velocity dispersion of 212 , which represents the median dispersion of the three elliptical galaxies in the sample. Although the grid is essential for estimating light-weighted stellar population properties, we use it here solely to validate our pPXF results, in order to verify their reliability in the presence of potential intrinsic uncertainties.
@@ 3.4 Voronoi binning for spatially resolved properties
As a complementary analysis, the stellar populations across the spaxels of the datacube were measured and visualised as a 2D map to study the properties of the inner and outer regions of the galaxies. Given the variations in S/N across the spaxels in the MUSE datacube, we opted to employ Voronoi binning which ensures that a minimum user-defined S/N per bin is achieved throughout the galaxy. The size of the Voronoi bins adapt based on the corresponding S/N, which allows for the spaxels of low S/N to be handled effectively (Cappellari & Copin, 2003) by assigning them to larger bins, while high S/N spaxels, usually located in central galaxy regions, are assigned to smaller bins or left unchanged (Cappellari & Copin, 2003).
The Python implementation of the Vorbin package of Cappellari & Copin (2003) was used to Voronoi bin the S/N in our galaxy datacubes. The continuum S/N was determined at a rest-frame wavelength of 5635 Å, in a region without any spectral features. According to Zibetti et al. (2024), a minimum spectral S/N of 20 is required to constrain stellar ages reliably from stellar population synthesis (SPS) models (see also Cid Fernandes et al., 2005; Riffel et al., 2009). We therefore chose the constant S/N threshold to be more conservative at 50 , following Johnston et al. (2021) where it was proven sufficient to extract stellar populations. The binning is performed such that the total S/N of each assigned bin is above the set threshold, thereby optimising the usage of spaxels on the outskirts of the galaxy with low S/N in addition to the spaxels with high S/N in the core of the galaxy. Any foreground stars in the MUSE field of view or neighbouring objects were masked out wherever necessary, as well as the spaxels with S/N to discard noise-dominated spaxels and sky spaxels.
The spectra in each bin were then summed, and modelled with pPXF, following the prescription detailed in Sect. 3.2 in order to extract the mean stellar metallicities and stellar ages. We note that this technique provides only a general overview of the stellar population properties of the galaxy in a 2D map, and allows us to study these properties and any features in “inner” and “outer” regions of the galaxy. While this is useful as a confirmation analysis, it is not an exact alternative to IFU galaxy decomposition described in Sect. 3.1. The Voronoi binned maps represent the light of the different components superposed on each other as the total light in each spaxel, and therefore propagates in each bin. BUDDI on the other hand extracts the light from each underlying component of the galaxy, minimising the superposition.
@@ 4.1 Component models and structural parameters
As described in Sect. 3.1, the galaxies were modelled with a single Sérsic profile at first, and then the complexity of the models were subsequently increased and tested using BUDDI in order to isolate any relevant substructure in elliptical galaxies presenting different physical and stellar population properties. We attempted to go as far as three components, including a double Sérsic profile with a PSF model, and a triple Sérsic component model, drawing from the findings in Huang et al. (2013a) and Huang et al. (2013b). However, adding a third component caused the fit to crash before completion due to a forced attempt at fitting a non-existent component for two of the galaxies (J205050 and J225546). Adding a PSF component to the double Sérsic model did not cause the fits to fail, but did not contribute to a significant difference to the models and spectra. For J020536, the fit appeared to be successful with three Sérsic components, only considering the residual image and the estimated structural parameters. However, on closer inspection, the fit had in fact resulted in spectra with a distorted and erratic continuum, suggesting inaccuracies due to significant noise. Given the spatial resolution offered by MUSE and the sub-optimal observing conditions, a double Sérsic model was chosen to be the best fit for all three galaxies, which was decided after visual inspection of the fit residuals and the estimated structural parameters, along with the extracted spectra. The structural parameters estimated by BUDDI are listed in Table 7 and the fits to the galaxies in Figure 1.
@@
Considering the two-component fits, the Sérsic indices of all the ellipticals were found to be less than 2 for both components, and therefore are not represented by de Vaucouleurs profiles. For J205050 and J225546, the components show , while J020536 shows the opposite. The axis ratios of the inner component is higher than the outer component in J020536 and J225546, while J205050 has nearly equal axis ratios for both components. For the single Sérsic fit, the Sérsic index is significantly higher () than either of the components of the double Sérsic fit. This higher Sérsic index reflects the attempt to fit both the steep and shallow light profiles dominating at different scales in the galaxy simultaneously. For a galaxy that has more than one component, this often results in an average value that does not represent either component accurately, but rather provides a compromise whilst fitting the entire galaxy, leading to the impression of a highly compact object. Similarly, this affects the sizes of the galaxies, with a single Sérsic fit yielding an that falls between the values of the two components.
@@
As is evident from the white-light images of the three ellipticals in Figure 1, there are a number of foreground stars or neighbouring galaxies within the field of view, which have the potential to affect the component models if not accounted for. For example, another object can be clearly observed close to J020536. Lassen et al. (2021) concluded that this object was an irregular dwarf galaxy at that is interacting with J020536, creating a plume of irregular structure between both objects. The asymmetric distortion in the surface brightness caused by the interaction makes the dwarf galaxy difficult to model simultaneously with the target. To avoid the fit from crashing or modelling the wrong object, we mask this entire region out with a polygon, in addition to masking out the stars. However, we note that despite our efforts to mask out the interacting object, the overlap between the galaxies implies that the light contamination from the dwarf galaxy in J020536 cannot be entirely eliminated, especially at the outskirts. Similarly, there is a small galaxy that appears close in projection to J225546. However, with a redshift of 0.22, it is clear that this object is neither a satellite galaxy accreting onto the target nor interacting with it. It is rather a background galaxy along the same line of sight as the main galaxy. However, since the projection of this object appears to lie very much within the outer regions of the target galaxy itself, it was decided to model this as an additional component. Since the “neighbouring” background galaxy is an extended object, discerning the point where its light contribution reaches negligible levels becomes challenging, making it difficult to mask. Choosing the size of the mask can have a significant effect in modelling the light from the main galaxy. In this instance, the mask could either be inadequate to allow light from the outskirts to influence the fit to the target galaxy, or excessively large, resulting in substantial loss of light from the main galaxy. For these reasons, we elected to model the background galaxy, which helps to subtract its light contribution, providing a more reliable fit to the galaxy of interest as compared to masking it out (Häussler et al., 2007).
@@
The two-component models of the galaxies as depicted in Figure 1 show a good fit, barring the very innermost regions. It must be noted that the seeing is indeed accounted for in these fits by convolving the flux datacubes with their PSF datacubes, but constructing an accurate PSF from IFU data presents a significant challenge, and can often prove hard to model the central few pixels. The lower right panel depicts the fraction of light that has been over- or under-subtracted from the white-light image as a function of the galactocentric distance along the major axis. The majority of the residuals lie within of the observed points for up to 20′′ from the galactic centre. The higher scatter seen at larger distances from the centre are likely due to the asymmetric sky background and the effect of low S/N in the outskirts. The galaxy 1D profile was constructed by plotting the surface brightness from the centre of the observed galaxy to the outskirts using a simulated “slit” along the major axis, following the PA derived from BUDDI for the single Sérsic model. The model profiles were extracted with the same slit along the major axis, accounting for the PSF smearing at the centre in a consistent way for the inner and outer components and the combined model datacubes, which had all been convolved with the PSF in the initial steps of 2D fitting. The 1D and 2D surface brightness profiles can be considered a 1:1 correspondence, constructed from the white-light images collapsed along the wavelength dimension. While this serves to support the reliable quality of the fits to the surface brightness of the galaxies, it is important to note that the 1D profiles rarely capture mismatches with the surface brightness models (depending on the choice of sampling rate and that are clearly visible spatially in the 2D profiles.

### v2

@@ 3.1 Galaxy decomposition with BUDDI
Having modelled the surface brightness of the galaxies with one, two, and three Sérsic profiles, we opted for the two-component models for all of them, which are shown in Fig. 1. To unveil the physical properties and implications associated with the different components of elliptical galaxies, the clean spectrum of each component was extracted with BUDDI. The magnitudes from GalfitM correspond to the total fluxes of each Sérsic profile integrated out to infinity. This extends to the individual image slice models in the datacube, and is reflected in the final 1D spectrum derived by BUDDI, which was obtained by plotting these fluxes as a function of wavelength (Johnston et al., 2017). The decomposed spectra of the inner and outer components, the sky, and that corresponding to the best fit model (inner + outer + sky) are shown for J020536 in Fig. 2, and in Figs. 10 and 11 for J205050 and J225546, respectively. We note that since most of the observing nights for the filler programme included high lunar illumination (Sect. 2), we have included an arbitrarily scaled scattered lunar model spectrum for each galaxy, obtained from the ESO SkyCalc 444Available from https://www.eso.org/observing/etc/bin/gen/form?INS.MODE=swspectr+INS.NAME=SKYCALC Sky Model Calculator tool (Noll et al., 2012; Jones et al., 2013). The effect of scattered moonlight is strong for galaxies J205050 and J225546, and closely matches the slope of the sky spectrum modelled by BUDDI, effectively demonstrating its ability to cleanly disentangle the light from physical galaxy components and from sky contributions. For J020536, the scattered moonlight has less of a contribution to the modelled sky spectrum. However, since the MUSE datacubes undergo two sky subtraction steps (once within the EsoRex pipeline, and later enhanced with the ZAP code), the lunar contribution could in principle be partially eliminated, leading to the slope differences between the scattered moonlight spectrum and the sky spectrum, especially in J020536. An important point to note is that wherever there is a known telluric feature in the lunar spectrum (for instance, the O2 A-band at Å in absorption), there appears a spectral feature in the sky spectrum modelled by BUDDI. The spectral feature in the sky spectrum may or not be in absorption due to the uncertainties in sky subtraction for the MUSE cube (including possible over-subtraction in some cases); it is nevertheless encouraging that BUDDI picks up these features at the expected wavelength. Software is often only as effective as the quality of the data it processes - even so, it has clearly demonstrated its effectiveness in cleanly decomposing the component spectra from the sky spectrum. The decomposed spectra of the components were then used for the stellar population analyses, in order to unveil the physical properties and implications associated with the different components of elliptical galaxies.
@@ 3.2 Spectral fitting for stellar population analysis
The spectra of the two components obtained from the previous step (Sect. 3.1) were fitted using pPXF to derive the stellar metallicity [M/H] and stellar age. The Python implementation of pPXF v9.1.1 was used in this work. pPXF finds a linear combination of stellar templates that best matches the input spectrum. For this study, both the mass-weighted and light-weighted stellar population properties were derived and analysed. The full spectral fitting method optimises the mass or light weights for each SSP spectrum in the template library, such that the observed spectrum is a combination of these weights (Cappellari & Copin, 2003; Cappellari, 2017, 2023). The mean logarithmic mass and light-weighted stellar ages and metallicities estimated by pPXF are defined as:
@@
The empirical library of stellar spectra employed was MILES (Vazdekis et al., 2015)555Available from http://miles.iac.es/pages/webtools.php, along with the BaSTI isochrone models (Pietrinferni et al., 2004), which track stellar evolution. This choice of the stellar library is attributed to the vast coverage of stellar metallicities and ages, which is necessary for studying nearby galaxies in the local Universe, which can contain both old and young stars with low and high metallicities. Restricting to the safe ranges defined in Vazdekis et al. (2010), the chosen stellar population parameters span a range from -1.79 to 0.26 dex for the metallicities [M/H] and from 30 Myr to 14 Gyr for the ages, forming a combined total of 530 template spectra. Since each stellar template is normalised to 1, pPXF derives the mass weights by default. In order to estimate the light weights, the stellar templates were normalised by the band luminosity ( Å), chosen such that the resulting properties favour more recent star formation episodes in the galaxy. The Universal Kroupa IMF (Kroupa & Boily, 2002) with a slope of 1.3 is assumed.
@@
During the spectral fit with pPXF, an 8th-order multiplicative Legendre polynomial was used to correct for the shape of the continuum. Furthermore, the inclusion of a multiplicative polynomial in the fit also ensures that the spectral fit is independent of dust reddening, thereby eliminating the need to define a reddening curve explicitly within the pPXF routine (Cappellari, 2017). Additionally, the [OI]5577Å sky line was masked during the fit. The code allows for emission lines to be fitted simultaneously alongside the stellar continuum using a set of respective Gaussian templates representing the nebular emission. The spectral range for the fit was constrained to rest-frame Å due to inconsistencies that can be introduced when using the full MUSE wavelength range (Krajnović et al., 2015). Moreover, this chosen range spans the important spectral features that are sensitive to stellar age and metallicity (, , , and Fe) and aligns with the wavelength coverage in the MILES models.
@@
The pPXF routine provides a solution to the ill-posed problem of extracting stellar population properties from observed spectra. The extensive span of the parameter space of the SSPs inevitably leads to non-unique solutions that are degenerate in metallicity or age. This is one of the commonly cited challenges in the fossil-record method with stellar population modelling and full spectral fitting, along with the imprecise age resolution for older populations with Gyr (Cid Fernandes et al., 2005; Riffel et al., 2009; Ibarra-Medel et al., 2019; Lacerna et al., 2020; Riffel et al., 2024). The solution incorporated within pPXF is regularisation, which is responsible for mitigating intrinsic degeneracies in the final solution. When a possible degeneracy is encountered, the inclusion of a regularisation term treats this by smearing out the weights of similar metallicities and ages, going from discrete weights that resemble a single stellar population to a solution with smooth weights that are more physically motivated in galaxies. The REGUL parameter within the routine sets the degree of smearing, following the prescription outlined in the pPXF user manual and in Shetty & Cappellari (2015) and Jegatheesan et al. (2024). A word of caution is advised at this point while interpreting the SFHs from full spectral fitting: star-formation is a stochastic and unpredictable process, which can occur on timescales shorter than those specified in the synthetic template models. The spectral fits for each component spectrum in the three elliptical galaxies resulting from pPXF are displayed in Fig. 3. We find substantially robust fits for all objects, with the exception of the outer of component of J205050. From Fig. 10, this galaxy appears to have the strongest lunar contamination, which has mostly been modelled off as the sky spectrum. However, if there is a fraction of residual scattered moonlight contaminating especially the outer regions, this can add to the noise of the spectrum of the outer component. This could lead to a less robust fit to the spectrum using pPXF; however, we still find that the crucial spectral features required for stellar population analysis have been modelled fairly well, and though we caution the direct interpretation of the fits and the resulting stellar populations for this component, we note that the SFHs are consistent with the other galaxies.
@@ 3.3 Spectral index measurements for stellar population analysis
We additionally measured the line strengths of stellar age and metallicity sensitive absorption features in the spectra through the Lick indices defined in Worthey (1999). This analysis was done as a consistency check to empirically determine the stellar populations from the spectra, due to potential model-dependent uncertainties and the effects of trade-offs achieved with the choice of the regularisation parameter, which can be introduced in the age-metallicity distribution estimated through full spectral fitting with pPXF. The absorption line strengths were measured using the public Python library PyLick666Available from https://gitlab.com/mmoresco/pylick/ (Borghi et al., 2022), which performs a fast estimation of the predominant Lick indices and their uncertainties, spanning the UV to the near-infrared region. The uncertainties were estimated following the method of Cardiel et al. (1998), through simulations accounting for the errors in the LOS velocities and S/N in the observed spectrum.
The H index served as the spectral diagnostic for stellar age, while the metallicity was determined using the composite index [MgFe]′ = , which is less sensitive to the -element abundance and serves as a more robust indicator of the total metallicity of the galaxy or component. These indices were plotted on a H[MgFe]′ plane in Fig. 5, overlaid with the SSP model grids defined in Vazdekis et al. (2010). The chosen grid spans ages from Gyr and metallicities [M/H] from -1.79 to +0.26 dex, constructed with the Lick index measurements from the MILES stellar library (Sánchez-Blázquez et al., 2006). The flexibility of the MILES webtools lies in its ability to customise the SSP models by convolving the model spectra with the velocity dispersion of the galaxy, effectively accounting for line broadening effects in the observed spectrum. As such, we used a convolved model with a velocity dispersion of 212 , which represents the median dispersion of the three elliptical galaxies in the sample. Although the grid is essential for estimating light-weighted stellar population properties, we have used it here solely to validate our pPXF results, in order to verify their reliability in the presence of potential intrinsic uncertainties.
@@ 3.4 Voronoi binning for spatially resolved properties
As a complementary analysis, the stellar populations across the spaxels of the datacube were measured and visualised as a 2D map to study the properties of the inner and outer regions of the galaxies. Given the variations in S/N across the spaxels in the MUSE datacube, we opted to employ Voronoi binning, which ensures that a minimum user-defined S/N per bin is achieved throughout the galaxy. The size of the Voronoi bins adapts based on the corresponding S/N, which allows for the spaxels of low S/N to be handled effectively (Cappellari & Copin, 2003) by assigning them to larger bins, while high S/N spaxels, usually located in central galaxy regions, are assigned to smaller bins or left unchanged (Cappellari & Copin, 2003).
The Python implementation of the Vorbin package of Cappellari & Copin (2003) was used to Voronoi bin the S/N in our galaxy datacubes. The continuum S/N was determined at a rest-frame wavelength of 5635 Å, in a region without any spectral features. According to Zibetti et al. (2024), a minimum spectral S/N of 20 is required to constrain stellar ages reliably from stellar population synthesis (SPS) models (see also Cid Fernandes et al., 2005; Riffel et al., 2009). We therefore chose the constant S/N threshold to be more conservative at 50 , following Johnston et al. (2021), in which it was proven to be sufficient to extract stellar populations. The binning was performed such that the total S/N of each assigned bin was above the set threshold, thereby optimising the usage of spaxels on the outskirts of the galaxy with low S/N in addition to the spaxels with high S/N in the core of the galaxy. Any foreground stars in the MUSE field of view or neighbouring objects were masked out wherever necessary, as well as the spaxels with S/N to discard noise-dominated spaxels and sky spaxels.
The spectra in each bin were then summed, and modelled with pPXF, following the prescription detailed in Sect. 3.2, in order to extract the mean stellar metallicities and stellar ages. We note that this technique provides only a general overview of the stellar population properties of the galaxy in a 2D map, and allows us to study these properties and any features in inner and outer regions of the galaxy. While this is useful as a confirmation analysis, it is not an exact alternative to the IFU galaxy decomposition described in Sect. 3.1. The Voronoi-binned maps represent the light of the different components superposed on each other as the total light in each spaxel, and therefore this contamination propagates in each bin. BUDDI, on the other hand, extracts the light from each underlying component of the galaxy, minimising the superposition.
@@ 4.1 Component models and structural parameters
As is described in Sect. 3.1, the galaxies were modelled with a single Sérsic profile at first, and then the complexity of the models was subsequently increased and tested using BUDDI in order to isolate any relevant substructure in elliptical galaxies presenting different physical and stellar population properties. We attempted to go as far as three components, including a double Sérsic profile with a PSF model, and a triple Sérsic component model, drawing from the findings in Huang et al. (2013a) and Huang et al. (2013b). However, adding a third component caused the fit to crash before completion due to a forced attempt at fitting a non-existent component for two of the galaxies (J205050 and J225546). Adding a PSF component to the double Sérsic model did not cause the fits to fail, but did not contribute to a significant difference to the models and spectra. For J020536, the fit appeared to be successful with three Sérsic components, only considering the residual image and the estimated structural parameters. However, on closer inspection, the fit had in fact resulted in spectra with a distorted and erratic continuum, suggesting inaccuracies due to significant noise. Given the spatial resolution offered by MUSE and the sub-optimal observing conditions, a double Sérsic model was chosen to be the best fit for all three galaxies, which was decided after visual inspection of the fit residuals and the estimated structural parameters, along with the extracted spectra. The structural parameters estimated by BUDDI are listed in Table 7 and the fits to the galaxies in Fig. 1.
@@
Considering the two-component fits, the Sérsic indices of all the ellipticals were found to be less than 2 for both components, and therefore are not represented by de Vaucouleurs profiles. For J205050 and J225546, the components show , while J020536 shows the opposite. The axis ratio of the inner component is higher than the outer component in J020536 and J225546, while J205050 has nearly equal axis ratios for both components. For the single Sérsic fit, the Sérsic index is significantly higher () than either of the components of the double Sérsic fit. This higher Sérsic index reflects the attempt to fit both the steep and shallow light profiles dominating at different scales in the galaxy simultaneously. For a galaxy that has more than one component, this often results in an average value that does not represent either component accurately, but that rather provides a compromise whilst fitting the entire galaxy, leading to the impression of a highly compact object. Similarly, this affects the sizes of the galaxies, with a single Sérsic fit yielding an that falls between the values of the two components.
@@
As is evident from the white-light images of the three ellipticals in Fig. 1, there are a number of foreground stars or neighbouring galaxies within the field of view, which have the potential to affect the component models if not accounted for. For example, another object can be clearly observed close to J020536. Lassen et al. (2021) concluded that this object was an irregular dwarf galaxy at that is interacting with J020536, creating a plume of irregular structure between both objects. The asymmetric distortion in the surface brightness caused by the interaction makes the dwarf galaxy difficult to model simultaneously with the target. To avoid the fit from crashing or modelling the wrong object, we masked this entire region out with a polygon, in addition to masking out the stars. However, we note that despite our efforts to mask out the interacting object, the overlap between the galaxies implies that the light contamination from the dwarf galaxy in J020536 cannot be entirely eliminated, especially at the outskirts. Similarly, there is a small galaxy that appears close in projection to J225546. However, with a redshift of 0.22, it is clear that this object is neither a satellite galaxy accreting onto the target nor interacting with it. Rather, it is a background galaxy along the same line of sight as the main galaxy. However, since the projection of this object appears to lie very much within the outer regions of the target galaxy itself, it was decided to model this as an additional component. Since the ‘neighbouring’ background galaxy is an extended object, discerning the point where its light contribution reaches negligible levels becomes challenging, making it difficult to mask. Choosing the size of the mask can have a significant effect in modelling the light from the main galaxy. In this instance, the mask could either be inadequate, allowing light from the outskirts to influence the fit to the target galaxy, or excessively large, resulting in substantial loss of light from the main galaxy. For these reasons, we elected to model the background galaxy, which helped us to subtract its light contribution, providing a more reliable fit to the galaxy of interest than masking it out would (Häussler et al., 2007).
@@
The two-component models of the galaxies depicted in Fig. 1 show a good fit, barring the innermost regions. It must be noted that the seeing is indeed accounted for in these fits by convolving the flux datacubes with their PSF datacubes, but constructing an accurate PSF from IFU data presents a significant challenge, and it can often prove hard to model the central few pixels. The lower right panel depicts the fraction of light that has been over- or under-subtracted from the white-light image as a function of the galactocentric distance along the major axis. The majority of the residuals lie within of the observed points for up to 20′′ from the galactic centre. The higher scatter seen at larger distances from the centre are likely due to the asymmetric sky background and the effect of low S/N in the outskirts. The galaxy 1D profile was constructed by plotting the surface brightness from the centre of the observed galaxy to the outskirts using a simulated ‘slit’ along the major axis, following the PA derived from BUDDI for the single Sérsic model. The model profiles were extracted with the same slit along the major axis, accounting for the PSF smearing at the centre in a consistent way for the inner and outer components and the combined model datacubes, which had all been convolved with the PSF in the initial steps of 2D fitting. The 1D and 2D surface brightness profiles can be considered a 1:1 correspondence, constructed from the white-light images collapsed along the wavelength dimension. While this serves to support the reliable quality of the fits to the surface brightness of the galaxies, it is important to note that the 1D profiles rarely capture mismatches with the surface brightness models (depending on the choice of sampling rate and that are clearly visible spatially in the 2D profiles.